# Learned harmonic mean estimation of the Bayesian evidence with normalizing flows

### v1

@@ 1 Introduction
In practice, the computation of Bayesian evidence is very challenging as it involves evaluating a multi-dimensional integral over a potentially highly varied function. The most widespread method for estimating the Bayesian evidence, particularly in astrophysics, is nested sampling (Skilling, 2006; Ashton et al., 2022; Buchner, 2021). While nested sampling has been highly successful and many effective nested sampling algorithms and codes have been developed (Feroz & Hobson, 2008; Feroz et al., 2009a, b; Brewer et al., 2011; Handley et al., 2015a, b; Feroz et al., 2019; Speagle, 2020; Buchner, 2021; Williams et al., 2021; Cai et al., 2022), it imposes constraints on the method used to sample. By sampling in a nested manner it is possible to reparameterize the likelihood in terms of the enclosed prior volume such that the evidence can be computed by a one-dimensional integral. The computational challenge then shifts to how to effectively sample in a nested manner, i.e. how to sample from the prior subject to likelihood level-sets or isocontours. The need to sample in this nested manner severely reduces flexibility (hence the need to design custom nested sampling algorithms), typically restricting application to relatively low dimensional settings. [^1] A notable exception that is applicable to high-dimensional settings is proximal nested sampling (Cai et al., 2022), although it is only applicable for log-convex likelihoods.
The harmonic mean estimator of the Bayesian evidence, introduced by Newton & Raftery (1994), provides much greater flexibility since it only requires samples from the posterior, available from any Markov chain Monte Carlo (MCMC) method, for example. However, it was immediately realized by Neal (1994) that the method can easily fail catastrophically due to the estimator’s variance becoming very large. To solve this issue the learned harmonic mean estimator was recently proposed by McEwen et al. (2021), where machine learning techniques were developed to learn a suitable internal importance sampling target distribution. Since the estimator requires only samples from the posterior and so is agnostic to the method used to generate samples, in contrast to nested sampling, it can be easily applied with any MCMC sampling technique, including saved down MCMC chains, or any variational inference approach. This property also allows the estimator to be adapted to address Bayesian model selection for simulation-based inference (SBI) (Spurio Mancini et al., 2023), where an explicit likelihood is unavailable or infeasible.
@@
While normalizing flows can learn a normalized posterior density by definition, the normalization constant itself, i.e. the Bayesian evidence, is not directly accessible. Nevertheless, the Bayesian evidence can be computed by backing out the normalization constant, as discussed in Spurio Mancini et al. (2023), by taking the ratio of the unnormalized posterior (given by the product of the likelihood and prior) with the normalizing flow representing a surrogate for the posterior. This approach, which we call the naïve normalizing flow estimator in Spurio Mancini et al. (2023), is highly dependent on the accuracy of the approximating normalizing flow and suffers a large variance, as discussed in Spurio Mancini et al. (2023). For comparison, we compute this naïve estimator in the current article and demonstrate its large variance. Very recently, Srinivasan et al. (2024) adopt this naïve estimator and attempt to reduce its variance by introducing an additional term in the loss that penalizes variability when training the flow. While this reduces the variability of the estimator, the estimator nevertheless remains highly dependent on the accuracy of the approximating flow and there is no guarantee that resulting evidence estimates are unbiased. Training flows using the forward Kullback-Leibler (KL) divergence when given samples from the target distribution is known to suffer from mode seeking behaviour, where the learned flow has narrower tails than the target (e.g. Murphy, 2022). While the approach presented in Srinivasan et al. (2024) suffers from this problem as it directly impacts the accuracy of estimated evidence values, our learned harmonic mean estimator does not. Firstly, in the learned harmonic mean approach the internal importance sampling target distribution that is learned should in any case have narrower tails than the posterior. Secondly, the distribution that is learned in our approach is not used as a surrogate for the posterior so it need not be an accurate approximation. For further details see Section 2.3 or McEwen et al., 2021.
@@ 3.2.1 Training the flow
Minimizing this approximation is equivalent to fitting the normalizing flow to the samples by maximum likelihood (Papamakarios et al., 2021). We take this approach to trainining our flow, and minimize the loss given by Equation (18) using the Adam optimizer (Kingma & Ba, 2017; Dozat, 2016). We use a portion of the samples for training the flow and reserve the rest to be used for inference, to be substituted when estimating the evidence.
@@
As discussed in Section 1, training a flow by forward KL divergence, or equivalently maximum likelihood, can suffer from mode seeking behaviour, where the learned flow has narrower tails than the target (e.g. Murphy, 2022). Nevertheless, this is not a problem for use of the flow in the learned harmonic estimator since we seek a learned distribution that has narrower tails than the posterior. Moreover, the learned distribution is not used as a surrogate for the posterior but rather as the importance sampling target. The learned distribution need not be a highly accurate approximation of the posterior but it critically must be concentrated within the posterior. Consequently, the mode seeking behaviour of training normalizing flows by forward KL divergence does not compromise the learned harmonic mean estimator.
@@ 3.2.5 Code
The learned harmonic mean estimator with normalizing flows is implemented in the harmonic package [^3] https://github.com/astro-informatics/harmonic, from version 1.2.0 onwards. The methodology described in this section, with real NVP and rational quadratic spline flows, has been implemented in JAX and is available in recent releases of harmonic on PyPi and GitHub. Furthermore, other parts of the harmonic code have been updated to run on JAX, which is a Python package offering acceleration, just-in-time compilation and automatic differentiation functionality (Bradbury et al., 2018). Consequently, harmonic can now be run on hardware accelerators such as GPUs, potentially reducing computation times and allowing the user to tackle more complex, computationally demanding problems. Additionally, the automatic differentiation functionality opens up the possibility of optimizing based on evidence (e.g. for experimental design), as gradients are now accessible all the way down to evidence level, which provides an intriguing avenue for further research. The normalizing flow portion of the code is implemented using the flax (Heek et al., 2023), TensorFlow Probability (Dillon et al., 2017), optax and distrax (DeepMind et al., 2020) packages.
@@ 3.2.6 Naïve Bayesian evidence estimation using normalizing flows
While posterior samples are typically available and hence used, in principle the samples $\theta_{i}$ do not necessarily need to be drawn from the posterior. An overall estimate of the evidence and its spread can then simply be computed from the mean of these evidence estimates and their standard deviation. However, the resulting evidence estimator is likely to be biased and will have a large variance. This is because it is highly dependent on the accuracy of the learned normalizing flow, which is used as a surrogate for the posterior, as discussed in Section 1. The learned harmonic mean does not suffer from these issues.
@@ 4.1 Architectures, sampling and training
In the experiments where a real NVP flow is used, translation networks of the affine coupling layers are given by two-layer dense neural networks with a leaky ReLU activation function in between. For the scaling layers this is scaled by another dense layer with a softplus activation. We permute the inputs between coupling layers to ensure the flow transforms all elements. In the low dimensional benchmark experiments we consider a real NVP flow, unless otherwise stated, with six coupling layers, where typically only the first two include scaling. When we use a rational quadratic spline flow, it has a range $-10$ to $10$ (outside of this range it defaults to a linear transformation). The conditioner for the spline hyperparameters is a multi-layer perceptron with a tanh activation. We use a Gaussian base distribution with zero mean and an identity covariance matrix for all flows.
@@ 4.2.2 Normal-Gamma
with mean $\mu_{0}=0$, shape $a_{0}=10^{-3}$ and rate $b_{0}=10^{-3}$. The precision scale factor $\tau_{0}$ controls how diffuse the prior is. Friel & Wyse (2012) apply the original harmonic mean for this example and show that the evidence estimate does not vary with $\tau_{0}$, unlike the analytic ground truth value. We repeat this experiment, drawing $1,500$ samples for $200$ chains, with burn in of $500$ samples, yielding $1,000$ posterior samples per chain. We use a real NVP flow with $2$ scaled and $4$ unscaled layers at temperatures $T=0.9$ and $T=0.95$ to estimate the evidence.
@@
Figure 4b shows the relative accuracy of the evidence estimate computed using our method for a range of prior sizes. It can be seen that, unlike the original harmonic mean estimator, our method is accurate for a range of $\tau_{0}$. Results are computed with the trained flow at temperature $T=0.9$ (blue) and $T=0.95$ (green). They are accurate in both cases, showing that the temperature parameter does not require fine-tuning.
@@ 4.2.3 Logistic regression models: Pima Indian example
This analysis was originally performed to study indicators of diabetes in $n=532$ Pima Indian women. The predictors of diabetes considered included the number of prior pregnancies (NP), plasma glucose concentration (PGC), body mass index (BMI), diabetes pedigree function (DP) and age (AGE). The probability of diabetes $p_{i}$ for person $i\in\{1,\ldots,n\}$ is modelled by the logistic function
@@
where both additionally include a bias. We estimate the Bayes factor of these models $\text{BF}_{12}$ with our method and compare it to a benchmark value computed by Friel & Wyse (2012) using a reversible jump algorithm (Green, 1995). They obtain a value of $\text{BF}_{12}=13.96$ ($\log\text{BF}_{12}=2.636$), which we treat as ground truth. We draw $5,000$ samples for $200$ chains, with burn in of 1,001001,001 , 000 samples, yielding $4,000$ posterior samples per chain. We use a real NVP flow with $2$ scaled and $4$ unscaled layers at temperature $T=0.9$, applying standardization.
@@ 4.2.4 Non-nested linear regression models: Radiata pine example
Using emcee, we draw $10,000$ samples for $200$ chains, with burn in of $2,000$ samples, yielding $8,000$ posterior samples per chain. We train a rational quadratic flow consisting of $2$ layers, with $50$ spline bins. Standardization is applied to the data as detailed in Section 3.2, which is necessary due to the vast difference in scale of the parameter dimensions.
@@ 4.3 Robustness of the temperature parameter
Method $\log(z_{\Lambda\text{CDM}})$ $\log(z_{w\text{CDM}})$ $\log\text{BF}_{\Lambda\text{CDM-$w$CDM}}$ Computation time (64 CPU cores)
@@
Many methods of estimating the evidence require careful fine-tuning of hyperparameters. As explained in Section 2.3, this was also the case for the learned harmonic mean estimator when using the classical machine learning models as considered previously. The target distribution was learned using simple machine learning models and a bespoke optimization approach designed to ensure the target distribution is contained within the posterior. The models had to be carefully chosen and their hyperparameters had to be fine-tuned for the estimator to be accurate and validated by cross-validation. In this work, through the introduction of a more sophisticated machine learning model, normalizing flows, we are able to avoid this drawback and create a more robust estimator.
@@ 4.4 Practical cosmological example: DES Y1 analysis
Note that the values computed by the learned harmonic mean and nested sampling are in close agreement, showing a slight preference for $\Lambda$CDM, matching the configuration of our simulated setup. The values computed by the naïve estimator are in approximate agreement but exhibit an error two orders or magnitude larger than the error of the learned harmonic mean (in higher dimensional examples to be reported in an ongoing work we observe the naïve estimator failing much more catastrophically).
@@ 5 Conclusions
Since the learned harmonic mean estimator is decoupled from the sampling method, it can be used in a wide variety of settings. This includes approaches such as simulation-based inference, variational inference and various MCMC methods where the evidence could not otherwise be computed accurately, such as the No U-Turn Sampler (NUTS) (Hoffman & Gelman, 2011). When using MCMC methods for parameter estimation, the Bayesian evidence can be obtained essentially “for free” or even post-hoc from saved down MCMC chains. Since the estimator is agnostic to the sampling strategy, it is highly flexible. The best suited sampling strategy may be used for the problem at hand, as we demonstrated in the DES Y1 example, where Metropolis-Hastings accurately sampled the posterior much faster than nested sampling. In ongoing work we leverage the flexibility of the learned harmonic mean to demonstrate its use with NUTS and the CosmoPower-JAX emulator (Spurio Mancini et al., 2022; Piras & Spurio Mancini, 2023) to scale evidence calculation to $\sim 150$ dimensions. Overall, the learned harmonic mean estimator with normalizing flows is a robust, flexible and scalable tool for Bayesian model comparison that can be used in a variety of contexts.

### v2

@@ 1 Introduction
In practice, the computation of Bayesian evidence is very challenging as it involves evaluating a multi-dimensional integral over a potentially highly varied function. The most widespread method for estimating the Bayesian evidence, particularly in astrophysics, is nested sampling (Skilling, 2006; Ashton et al., 2022; Buchner, 2021). While nested sampling has been highly successful and many effective nested sampling algorithms and codes have been developed (Feroz & Hobson, 2008; Feroz et al., 2009a, b; Brewer et al., 2011; Handley et al., 2015a, b; Feroz et al., 2019; Speagle, 2020; Buchner, 2021; Williams et al., 2021; Cai et al., 2022), it imposes constraints on the method used to sample. By sampling in a nested manner it is possible to reparameterize the likelihood in terms of the enclosed prior volume such that the evidence can be computed by a one-dimensional integral. The computational challenge then shifts to how to effectively sample in a nested manner, i.e. how to sample from the prior subject to likelihood level-sets or isocontours. The need to sample in this nested manner severely reduces flexibility (hence the need to design custom nested sampling algorithms), typically restricting application to relatively low-dimensional settings. [^1] A notable exception that is applicable to high-dimensional settings is proximal nested sampling (Cai et al., 2022), although it is only applicable for log-convex likelihoods.
The harmonic mean estimator of the Bayesian evidence, introduced by Newton & Raftery (1994), provides much greater flexibility since it only requires samples from the posterior, available from any Markov chain Monte Carlo (MCMC) method, for example. However, it was immediately realized by Neal (1994) that the method can easily fail catastrophically due to the estimator’s variance becoming very large. To solve this issue the learned harmonic mean estimator was recently proposed by McEwen et al. (2021), where machine learning techniques were developed to learn a suitable internal importance sampling target distribution. Other evidence estimation methods decoupled from the evidence have been proposed recently (Heavens et al., 2017; Jia & Seljak, 2020; Srinivasan et al., 2024). Since the estimator requires only samples from the posterior and so is agnostic to the method used to generate samples, in contrast to nested sampling, it can be easily applied with any MCMC sampling technique, including saved down MCMC chains, or any variational inference approach. This property also allows the estimator to be adapted to address Bayesian model selection for simulation-based inference (SBI) (Spurio Mancini et al., 2023), where an explicit likelihood is unavailable or infeasible.
@@
While normalizing flows can learn a normalized posterior density by definition, the normalization constant itself, i.e. the Bayesian evidence, is not directly accessible. Nevertheless, the Bayesian evidence can be computed by backing out the normalization constant, as discussed in Spurio Mancini et al. (2023), by taking the ratio of the unnormalized posterior (given by the product of the likelihood and prior) with the normalizing flow representing a surrogate for the posterior. This approach, which we call the naïve normalizing flow estimator in Spurio Mancini et al. (2023), is highly dependent on the accuracy of the approximating normalizing flow and suffers a large variance, as discussed in Spurio Mancini et al. (2023). For comparison, we compute this naïve estimator in the current article and demonstrate its large variance. Very recently, Srinivasan et al. (2024) adopt this naïve estimator and attempt to reduce its variance by introducing an additional term in the loss that penalizes variability when training the flow. While this reduces the variability of the estimator, the estimator nevertheless remains highly dependent on the accuracy of the approximating flow and there is no statistical guarantee that resulting evidence estimates are unbiased. Training flows using the forward Kullback-Leibler (KL) divergence when given samples from the target distribution is known to suffer from mode covering behaviour, where the learned flow has wider tails than the target (e.g. Murphy, 2012). While the approach presented in Srinivasan et al. (2024) suffers from this problem which can directly impact the accuracy of estimated evidence values, our learned harmonic mean estimator does not. Firstly, in the learned harmonic mean approach with normalizing flow presented in this article, we concentrate the probability density of the internal importance sampling target distribution that is learned. Secondly, the distribution that is learned in our approach is in any case not used as a surrogate for the posterior so it need not be an accurate approximation. For further details see Section 2.3 or McEwen et al., 2021.
@@ 3.2.1 Training the flow
Minimizing this approximation is equivalent to fitting the normalizing flow to the samples by maximum likelihood (Papamakarios et al., 2021). We take this approach to training our flow, and minimize the loss given by Equation (18) using the Adam optimizer (Kingma & Ba, 2017; Dozat, 2016). We use a portion of the samples for training the flow and reserve the rest to be used for inference, to be substituted when estimating the evidence.
@@ 3.2.5 Code
The learned harmonic mean estimator with normalizing flows is implemented in the harmonic package [^3] https://github.com/astro-informatics/harmonic, from version 1.2.0 onwards. The methodology described in this section, with real NVP and rational quadratic spline flows, has been implemented in JAX and is available in recent releases of harmonic on PyPi and GitHub. Furthermore, other parts of the harmonic code have been updated use JAX, which is a Python framework offering acceleration, just-in-time compilation and automatic differentiation functionality (Bradbury et al., 2018). Consequently, harmonic can now be run on hardware accelerators such as GPUs, potentially reducing computation times and allowing the user to tackle more complex, computationally demanding problems. Additionally, the automatic differentiation functionality opens up the possibility of optimizing based on evidence (e.g. for experimental design), as gradients are now accessible all the way down to evidence level, which provides an intriguing avenue for further research. The normalizing flow portion of the code is implemented using the flax (Heek et al., 2023), TensorFlow Probability (Dillon et al., 2017), optax and distrax (DeepMind et al., 2020) packages.
@@ 3.2.6 Naïve Bayesian evidence estimation using normalizing flows
While posterior samples are typically available and hence used, in principle the samples $\theta_{i}$ do not necessarily need to be drawn from the posterior. An overall estimate of the evidence and its spread can then simply be computed from the mean of these evidence estimates and their standard deviation. However, the resulting evidence estimator is likely to be biased and will have a large variance.
@@ 4.1 Architectures, sampling and training
In the experiments where a real NVP flow is used, translation networks of the affine coupling layers are given by two-layer dense neural networks with a leaky ReLU activation function in between. For the scaling layers this is scaled by another dense layer with a softplus activation. We permute the inputs between coupling layers to ensure the flow transforms all elements. In the low dimensional benchmark experiments we consider a real NVP flow, unless otherwise stated, with six coupling layers, where typically only the first two include scaling. When we use a rational quadratic spline flow, it has a range $-10$ to $10$ (outside of this range it defaults to a linear transformation). The conditioner for the spline hyperparameters is a multi-layer perceptron with a hyperbolic tangent activation. We use a Gaussian base distribution with zero mean and an identity covariance matrix for all flows.
@@ 4.2.2 Normal-Gamma
with mean $\mu_{0}=0$, shape $a_{0}=10^{-3}$ and rate $b_{0}=10^{-3}$. The precision scale factor $\tau_{0}$ controls how diffuse the prior is. Friel & Wyse (2012) apply the original harmonic mean for this example and show that the evidence estimate does not vary with $\tau_{0}$, unlike the analytic ground truth value. We repeat this experiment, drawing $1,500$ samples for $200$ chains, with burn-in of $500$ samples, yielding $1,000$ posterior samples per chain. We use a real NVP flow with $2$ scaled and $4$ unscaled layers at temperatures $T=0.9$ and $T=0.95$ to estimate the evidence.
@@
Figure 4b shows the relative accuracy of the evidence estimate computed using our method for a range of prior sizes. It can be seen that, unlike the original harmonic mean estimator, our method is accurate for a range of $\tau_{0}$. Results are computed with the trained flow at temperature $T=0.9$ (blue) and $T=0.95$ (green). They are accurate in both cases, showing that the temperature parameter does not require fine-tuning. A detailed discussion of this point is included in Section 4.4.
@@ 4.2.3 Logistic regression models: Pima Indian example
This analysis was originally performed to study indicators of diabetes in $n=532$ Pima Native American women. The predictors of diabetes considered included the number of prior pregnancies (NP), plasma glucose concentration (PGC), body mass index (BMI), diabetes pedigree function (DP) and age (AGE). The probability of diabetes $p_{i}$ for person $i\in\{1,\ldots,n\}$ is modelled by the logistic function
@@
where both additionally include a bias. We estimate the Bayes factor of these models $\text{BF}_{12}$ with our method and compare it to a benchmark value computed by Friel & Wyse (2012) using a reversible jump algorithm (Green, 1995). They obtain a value of $\text{BF}_{12}=13.96$ ($\log\text{BF}_{12}=2.636$), which we treat as ground truth. We draw $5,000$ samples for $200$ chains, with burn-in of $1,000$ samples, yielding $4,000$ posterior samples per chain. We use a real NVP flow with $2$ scaled and $4$ unscaled layers at temperature $T=0.9$, applying standardization.
@@ 4.2.4 Non-nested linear regression models: Radiata pine example
Using emcee, we draw $10,000$ samples for $200$ chains, with burn-in of $2,000$ samples, yielding $8,000$ posterior samples per chain. We train a rational quadratic flow consisting of $2$ layers, with $50$ spline bins. Standardization is applied to the data as detailed in Section 3.2, which is necessary due to the vast difference in scale of the parameter dimensions.
@@
4.3 Gaussian in 21 dimensions
Figure 7: Violin plot of the log reciprocal Bayesian evidence computed by the learned harmonic mean estimator for a 21-dimensional Gaussian benchmark problem repeated $100$ times at $T=0.8$. The plot shows log reciprocal Bayesian evidence estimates across runs (measured) along with the one estimate and its error estimate (estimated). The ground truth is shown in red.
To further validate the method and error estimate in a moderate-dimensional context where the ground truth is available, we consider a simple $21$-dimensional Gaussian example. We choose this number of dimensions as it is the same as for the cosmological example we consider in Section 4.5. The diagonal elements are initialised as one plus Gaussian noise with zero mean and unit variance scaled by $0.1$. The off-diagonal elements adjacent to the diagonal are set to be the geometric mean of their adjacent diagonal elements scaled by $0.5$, with alternating signs. We draw $5,000$ MCMC samples for $80$ chains, with burn-in of $500$ samples, yielding $4,500$ posterior samples per chain. We use a rational quadratic flow consisting of $3$ layers, with $128$ spline bins with standardization, at temperature $T=0.8$. We train on half of the chains and use the other half for inference. We repeat this experiment $100$ times with different seeds.
Figure 7 shows the results of this experiment. The analysis is analogous to the 2-dimensional Rosenbrock experiment described in Section 4.2.1, but in log scale. It can be seen that the estimated value are in agreement with the ground truth, and the estimated and measured errors are similar.
@@ 4.4 Robustness of the temperature parameter
(a) Log reciprocal evidence for 10D Rosenbrock
(b) Learned harmonic mean log reciprocal evidence for varying temperature
Figure 9: Log reciprocal evidence values for the 10D Rosenbrock example. (a) Comparison of estimates obtained using different methods. Starred methods denote values obtained by Lange (2023). (b) Estimates obtained using the learned harmonic mean for a range of temperature values. It can be seen that the estimator is robust to this hyperparameter.
@@
Method $\log(z_{\Lambda\text{CDM}})$ $\log(z_{w\text{CDM}})$ $\log\text{BF}_{\Lambda\text{CDM-$w$CDM}}$ Computation time (64 CPU cores for sampling)
@@
Many methods of estimating the evidence require careful fine-tuning of hyperparameters. As explained in Section 2.3, this was also the case for the learned harmonic mean estimator when using the classical machine learning models as considered previously. In this work, through the introduction of a more sophisticated machine learning model, normalizing flows, we are able to avoid this drawback and create a more robust estimator.
@@
To investigate the impact of the temperature parameter for a non-Gaussian example of moderate dimensions, we study a 10-dimensional Rosenbrock example considered by Lange (2023). We consider a range of temperatures $T\in[0.5,0.8]$. We draw $10,000,000$ samples for $200$ chains, with burn-in of $404,800$ samples, yielding $9,595,200$ posterior samples per chain. Note that this is a challenging density to sample in higher dimensions, and emcee is not the optimal choice for this task. For this reason, Lange (2023) consider a much higher number of samples, which we reduce due to computational constraints. Additionally, the aim of this work is not to apply MCMC methods in these challenging settings, but to validate the robustness of the learned harmonic mean for a moderate dimensional non-Gaussian example, hence we choose not to focus our resources on this complicated task. For each temperature, we consider subsets of the sampled chains: we reserve the last 0.2% of the chains for training, while leaving the rest for inference. We thin the training set by a factor of $5$, and the inference set by a factor of $1,000$. For each subset, we increase the thinning starting index by $10$.
The results of this analysis are shown in Figure 9. Figure 9a shows the comparison of our log reciprocal evidence estimate at $T=0.8$, alongside the values quoted by Lange (2023). Our estimate is in broad agreement with many nested sampling methods, while being significantly more precise and using fewer samples. Figure 9b additionally shows the variation of the log reciprocal evidence estimates with temperature: it can be seen that these results are self-consistent, and hence our estimator is robust to this hyperparameter in this setting as well.
@@ 4.5 Practical cosmological example: DES Y1 analysis
Note that the values computed by the learned harmonic mean and nested sampling are in agreement, showing a slight preference for $\Lambda$CDM, matching the configuration of our simulated setup. The values computed by the naïve estimator are in approximate agreement but exhibit an error two orders or magnitude larger than the error of the learned harmonic mean (in higher dimensional examples to be reported in an ongoing work we observe the naïve estimator failing much more catastrophically).
@@ 5 Conclusions
Since the learned harmonic mean estimator is decoupled from the sampling method, it can be used in a wide variety of settings. This includes approaches such as simulation-based inference, variational inference and various MCMC methods where the evidence could not otherwise be computed accurately, such as the No U-Turn Sampler (NUTS) (Hoffman & Gelman, 2011). When using MCMC methods for parameter estimation, the Bayesian evidence can be obtained essentially “for free” or even post-hoc from saved down MCMC chains. Since the estimator is agnostic to the sampling strategy, it is highly flexible. The best suited sampling strategy may be used for the problem at hand, as we demonstrated in the DES Y1 example, where Metropolis-Hastings accurately sampled the posterior much faster than nested sampling. In recent work we leverage the flexibility of the learned harmonic mean to demonstrate its use with NUTS and the CosmoPower-JAX emulator (Spurio Mancini et al., 2022; Piras & Spurio Mancini, 2023) to scale evidence calculation to $\sim 150$ dimensions (Piras et al., 2024). Overall, the learned harmonic mean estimator with normalizing flows is a robust, flexible and scalable tool for Bayesian model comparison that can be used in a variety of contexts.